OpenStudy (anonymous):

Change this radical to an algebraic expression with fractional exponents. A) √b Type the expressions as radicals. B) q3/2 Perform the following computation with radicals. Simplify the answer. C) √6 . √8 D) ^4√5x . ^4√2x

OpenStudy (anonymous):

\[\sqrt{b} = b ^{(1/2)}\]

OpenStudy (anonymous):

n^(a/b). The b is the power, and a is the exponent under the power . \[q ^{3/2} = \sqrt{q ^{3}}\]

OpenStudy (anonymous):

\[\sqrt{6} \times \sqrt{8} = \sqrt{6 \times 8} = \sqrt{48} = 4\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt[4]{5x} \times \sqrt[4]{2x} = \sqrt[4]{5x*2x} =\sqrt[4]{10x ^{2}} = \sqrt[4]{10} \sqrt{x}\]